34.20 g of CaO is dissolved in 500.0 mL water. What is the concentration of the O2- ion?


1
Expert's answer
2021-04-28T02:46:34-0400

CaO + H2O --> Ca(OH)2

Molar Mass of CaO = 56

500.0 = 0.5 L

34.2/56 = 0.6482M

OH = 1.2964 × 2 = 2.5929

= 2/0.0025929 + 0.5

= 2/0.5025929

= 3.9794M/L